10.2.21-T Question Help * Find the regression equation, letting the first variable be the predictor (x variable. Using the listed actress/actor ages in various years, find the best predicted age of the Best Actor winner given that the age of the Best Actress winner that year is 44 years. ls the result within 5 years of the actual Best Actor winner, whose age was 43 years? Best Actress 29 32 28 64 30 35 46 28 61 23 44 56 Best Actor 45 38 37 44 53 50 60 48 37 56 43 35 Find the equation of the regression line. Round the constant to one decimal place as needed. Round the coefficient to three decimal places as needed.)
Answer #1
X Y X * Y X2
29 45 1305 841
32 38 1216 1024
28 37 1036 784
64 44 2816 4096
30 53 1590 900
35 50 1750 1225
46 60 2760 2116
28 48 1344 784
61 37 2257 3721
23 56 1288 529
44 43 1892 1936
56 35 1960 3136
Total 476 546 21214 21092

\hat{Y} = a + bX

Where , a :- is the Y intercept

b :- is the slope of the equation

b = ( n \Sigma XY - \Sigma X \Sigma Y) / ( n\Sigma X^{2} - (\Sigma X)^{2})

b = ( 12 * 21214 - 476 * 546) / ( 12 * 21092 - (476)^{2})

b = -0.201

a =( \Sigma Y - ( b * \Sigma X) ) / n

a =( 546 - ( -0.201 * 476) ) / 12

a =53.5

\hat{Y} = 53.5 - 0.201 X

To find the best predicted age of the actor given the age of the actress is 44 years

\hat{Y} = 53.5 - 0.201 (44)

\hat{Y} =45

Actor's age is 45 years.